OpenStudy (anonymous):

angle 1and angle 2 are supplementary angles. angle 1 is 4y + 7 and angle 2 is 9y + 4. Find angle 2.

OpenStudy (anonymous):

13y+11=180 13y=169 y=13 second angle 9(13)+4 121

OpenStudy (aroub):

supplementary angles add up to 180 <1 +<2 =180 , now just substitute 4y+7+9y+4=180 13y+11=180 13y=169 y=13 now we found y ,so you substitute 13 by the y in angle 2 , so you could get the measure of angle 2 9y + 4 9(13)+4 = 121

OpenStudy (anonymous):

Supplementary means sum is 180 degrees or 2 \pi radians. So here, \[4y + 7 + 9y + 4 = 180\] \[\Rightarrow 13y+11=180\] \[13y=169\] \[y=13\] So,substituting for the second angle we get, \[ 9 \times (13)+4 = 121 \]

OpenStudy (anonymous):

hey lagrangeson,wats up with the madels? hm?(sarcastic)